Wesley Power Tools manufactures a wide variety of tools and accessories. One of its more popular items is a cordless power handisaw. Each handisaw sells for $70. Wesley expects the following unit sales

January 2,400
February 2,700
March 2,700
April 2,700
May 2,200

Wesley’s ending finished goods inventory policy is 35 percent of the next month’s sales.
      Suppose each handisaw takes approximately .55 hours to manufacture, and Wesley pays an average labor wage of $12.50 per hour.
      Each handisaw requires a plastic housing that Wesley purchases from a supplier at a cost of $7.00 each. The company has an ending raw materials inventory policy of 10 percent of the following month’s production requirements. Materials other than the housing unit total $4.50 per handisaw.
      Manufacturing overhead for this product includes $69,000 annual fixed overhead (based on production of 24,000 units) and $.90 per unit variable manufacturing overhead. Wesley’s selling expenses are 7 percent of sales dollars, and administrative expenses are fixed at $16,000 per month.

Required:
1.
Compute the budgeted cost of goods sold for the first quarter. (Round direct material, direct labor and overhead costs per unit to 2 decimal places. Round final answers to the nearest dollar amount.)

       

2. Compute the budgeted selling and administrative expenses.



3. Complete the budgeted income statement for the handisaw product for the first quarter. (Round direct material, direct labor and overhead costs per unit to 2 decimal places. Round final answers to the nearest dollar amount.)
Answer #1

1)Standard cost : Direct material +Direct labor +variable overhead+ fixed overhead

           = [(1*7) +4.5 ]+ (.55*12.5] + .90+ [69000/24000]

           = 7 + 4.5 + 6.88 + .90 + 2.88

           = $ 22.16 per unit

Total number of units sold in First quarter : 2400+2700+2700= 7800 Units

the budgeted cost of goods sold for the first quarter : Cost per unit * Units sold

                       = 22.16 *7800

                       = $ 172848

2)

Sales for first quarter (7800*70) 546000
Variable selling expense (546000* 7%) 38220
Fixed v administrative expenses (16000*3) 48000
Total budgeted selling and administrative expenses 86220

3)

WESLEY POWER TOOLS

INCOME STATEMENT

sales 546000
less:cost of goods sold (172848)
Gross margin 373152
less:selling and administrative expense (86220)
Net operating income 286932